OpenStudy (anonymous):

Find the magnitude and direction angle θ, to the nearest tenth of a degree, for the given vector v. v = -4i - 3

OpenStudy (anonymous):

I got 5; 216.9deg Nor real confident!!

OpenStudy (anonymous):

arctan(4/3)=53.1301023542

OpenStudy (jdoe0001):

-4i-3 standard form -3 - 4i ^ ^ a b components \(\large \bf || < -3, -4> || = \sqrt{(-3)^2+(-4)^2} \implies \sqrt{25} = 5\)

OpenStudy (jdoe0001):

$$\bf tan(\theta) = \cfrac{b}{a} \implies \frac{-4}{-3}\\ tan^{-1}(tan(\theta)) = tan^{-1}\left(\frac{-4}{-3}\right)\\ \theta = 53.13^o $$

OpenStudy (anonymous):

thanks to you both

OpenStudy (jdoe0001):

yw